# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v3

@@ Appendix C Representer theorem for RKHMs over general $C^{*}$-algebras
Hashimoto et al. (Hashimoto et al., 2023b, Proposition 4.5) showed an approximate representer theorem for RKHMs over von Neumann-algebras. Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
@@
Since $b_{i}\geq_{\mathcal{L}(\mathcal{M}_{k})}0$, there exists $c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$ such that $b_{i}=c^{2}$. Thus, $b_{i}-b_{i}^{2}=c^{2}-c^{4}=c(1-c^{2})c=c^{*}(1-b_{i})c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$. Thus, $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}^{2}\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1, and we have $|b_{i}f_{0}|_{k}^{2}=\left\langle f_{0},b_{i}^{2}f_{0}\right\rangle_{k}\leq_{\mathcal{A}}\left\langle f_{0},f_{0}\right\rangle_{k}=|f_{0}|_{k}$. As a result, we obtain
@@
Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition. $\Box$
Appendix D Generalization to Other $C^{*}$-algebras
Continuous functions on high-dimensional torus (Leimbach & van Suijlekom, 2024)
[…]
Continuous functions on the sphere (Rieffel, 2004)
Appendix E Experimental Details and Additional Results
We provide experimental details below. All the experiments in this paper were executed with Python 3.9 on an Intel(R) Core(TM) i9-10885H 2.4GHz processor with the Windows 10 operating system.
Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the regression task. (Average value of results of five different runs. The error bar represents the standard deviation.)
E.1 Experiment with synthetic data
We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$. We generated 1000 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimation of $f$ and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. For $x^{i}$, $y^{i}$, $\tilde{x}^{i}$, and $a$, we discretized each function with 30 equally spaced points on $\mathbb{T}$. In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$ discussed in Remark 15, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 3 shows the result. The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
We also show additional results regarding the deep approach in Figure 4. The result is the average $\pm$ the standard deviation of three independent runs. We set different values of $n$ with the same deep setting considered in Subsection 7.1.
Figure 4: Test error of the regression task with deep approach with different $n$ and $L$.
E.2 Experiment with MNIST
We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. Figure 5 (a) shows the test error. We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$. In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 5 (b) shows the result. The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.

### v4

@@ Appendix C Representer theorem for RKHMs over general $C^{*}$-algebras
Hashimoto et al. (Hashimoto et al., 2023b, Proposition 4.5) showed an approximate representer theorem for RKHMs over von Neumann-algebras. Here, von Neumann-algebras are $C^{*}$-algebras that is closed with respect to the strong operator topology. Since $C(\mathbb{T})$ is not a von Neumann-algebra, for the arguments in Section 5, we generalize the theorem to that for general $C^{*}$-algebras as follows.
@@
Since $b_{i}\geq_{\mathcal{L}(\mathcal{M}_{k})}0$, there exists $c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$ such that $b_{i}=c^{2}$. Thus, $b_{i}-b_{i}^{2}=c^{2}-c^{4}=c(1-c^{2})c=c^{*}(1-b_{i})c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$. Thus, $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}^{2}\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1, and we have $|b_{i}f_{0}|_{k}^{2}=\left\langle f_{0},b_{i}^{2}f_{0}\right\rangle_{k}\leq_{\mathcal{A}}\left\langle f_{0},f_{0}\right\rangle_{k}=|f_{0}|_{k}^{2}$. As a result, we obtain
@@
Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition.